OpenStudy (anonymous):

x-1/2 = x-2/3 << proportion x-1/2 = x-2/3 << proportion @Mathematics

OpenStudy (anonymous):

might this be \[\frac{x-1}{2}=\frac{x-2}{3}\]?

OpenStudy (anonymous):

if so "cross multiply" to get \[3(x-1)=2(x-2)\] \[3x-3=2x-4\] \[x-3=-4\] \[x=-1\]